OpenStudy (anonymous):

How many 3 digit positive integers N are there, such that N is a multiple of both 7 and 13?

OpenStudy (anonymous):

So N= _ _ _ where the spaces denote each digit of N N=7*13*1=91 so we move on to the next N=7*13*2=182 Here you found one N, which is 182 Try the next N=7*13*3=273 You found another N, and so on

OpenStudy (anonymous):

thanxx

OpenStudy (anonymous):

70 x 13= 910 only one for" just "

OpenStudy (anonymous):

Welcome :-)